P12-11B  

Prepare and post entries for partnership liquidation.

(LO 7) AP

The partners in Omni Services decided to liquidate the partnership on May 31, 2017, when balances in the company's accounts were as follows:

Item

Cash

Accounts Receivable

Equipment

Accumulated Depreciation

Accounts Payable

B. Hally, Capital

H. Lockyear, Capital

A. Vu, Capital

Balances before

liquidation

$33,000

$20,000

$75,200

$6,400

$53,160

$39,600

$25,200

$3,840

The partners share profit and loss 5:3:2 for Hally, Lockyear, and Vu, respectively.

Instructions

(a)  

Complete the schedule and prepare the journal entries for the liquidation of the partnership assuming the noncash assets were sold for $88,800, liabilities are paid, and the cash is distributed appropriately.
Answer #1

Particulars 317 May 6400 Accumulated Deprecation То белоруслф) 6400 Cash All To Accounts Recewable To Equipment 88800 Leova 2

WORKING NOTES :

value of equipment = 75200 - 6400 = 68800

value of cash = 33000+88800 = 121800

profits remained for the partner = 121800 - 53160(accounts payable) = 68640

share of hally = 68640*5/10 = 34320

share of lockyear = 68640*3/10 = 20592

share of vu = 68640*2/10 = 13728